OpenStudy (anonymous):

Simplify completely quantity 4 x plus 16 over 6. I took the quiz once before and i know that B is not the correct answer. The answer choices are: 2 open parentheses x plus 4 close parentheses over 3 2 open parentheses x plus 8 close parentheses over 3 4 open parentheses x plus 2 close parentheses over 3 4 open parentheses x plus 8 close parentheses over 3

OpenStudy (mathstudent55):

Is this the question? 

OpenStudy (mathstudent55):

If so, then start by factoring the numerator. The terms in the numerator are 4x and 16. What is the largest factor they have in common?

OpenStudy (anonymous):

yes & i dont really know how to factor.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the largest number that goes in to 4,16 and 6 evenly is just 2

OpenStudy (misty1212):

divide each term by 2

OpenStudy (misty1212):

\[\frac{4x+16}{6}=\frac{2x+3}{3}\] that is the best you can do

OpenStudy (anonymous):

i get what youre saying, thats not an answer choice though :(

OpenStudy (mathstudent55):

The largest factor of 4x and 16 is 4, so factor 4 out of the numerator. You get: